Servicing Mission Observatory Verification (SMOV) Update
As the first week of August comes to a close, most of Hubble’s science instruments have already completed or are close to completing their calibration activities. Each instrument has multiple channels that detect different wavelengths of light, and each channel must be tested and calibrated individually. While some instrument channels are still under evaluation, several others are already at work studying the universe.
- The Wide Field Camera 3 (WFC3) is nearing the end of its calibrations. Most calibrations for WFC3 should be completed next week. The camera has already completed enough of its calibration work to start taking images. It is continuing to take additional Early Release Observations (EROs), which will be shared with the public in about a month.
- Both the far-ultraviolet and near-ultraviolet channels of the Cosmic Origins Spectrograph (COS) continue to undergo calibrations as well. This week, scientists and engineers have been instructing COS to take measurements to check, among other things, the instrument’s sensitivity to light. Calibrations for COS will continue into September.
- The near-ultraviolet and far-ultraviolet channels of the Space Telescope Imaging Spectrograph (STIS) are also finishing up their calibration activities. The near-ultraviolet channel has shown higher-than-expected levels of internal ("dark") current, which engineers think might be due to the instrument being off since suffering a power supply failure five years ago this week. Engineers are making adjustments to accommodate for the extra current, and the adjustments appear to be working well so far. Calibrations will continue through next week.
Having quickly completed its calibrations, the STIS CCD channel took its first Early Release Observation at the end of June and will take another one later this month. The channel is routinely performing science operations.
- Following a failed attempt last week to restart the cooling system of the Near Infrared Camera and Multi-Object Spectrometer (NICMOS), engineers solved the problem and successfully restarted the cooling system on August 1. To their surprise and delight, NICMOS appears to be cooling more efficiently (faster) than expected.
Because NICMOS observes infrared wavelengths, which we detect as heat, the instrument has to be cold enough so that the heat of its own electronics doesn’t interfere with observations. It will take at least another week, possibly more, for NICMOS to reach operational temperatures of less than –321 degrees Fahrenheit (–196 degrees Celsius). Then the instrument team will have to wait a little while longer to make sure the temperature stabilizes sufficiently for science observations. At that point, engineers expect to be able to turn on NICMOS and begin its calibrations.
Space Telescope Science Institute (STScI)